# Improving Instruction-Following in Language Models through Activation Steering

### v1

@@ 5 Word-specific Instructions
Keyword Inclusion. For this set of instructions, we compute word-specific steering vectors. To generate these vectors, we append different phrasings of a request to include a specific word $w$ (e.g., “please include the word {$w$} in your response”) to a set of base queries. These prompts are used to compute a vector representation for the “include word {$w$}” instruction. While this requires a separate steering vector for each keyword at inference time, the vectors can be generated on the fly using arbitrary base queries unrelated to the keyword itself. Furthermore, these vectors can be computed with a small number of examples; in our experiments, we use only 20 examples. The models are then evaluated on the subset of IFEval that contains keyword inclusion/exclusion constraints. [^5] IFEval prompts may contain instructions for the inclusion or exclusion of multiple keywords in the same example. We separate such cases into inputs with single-keyword instructions.
@@
Keyword Exclusion. For keyword exclusion, we initially used the same procedure, appending instructions like “ensure the word $w$ does not appear in the answer” to base queries and computing the corresponding steering vectors. However, upon inspecting the steering vectors, we found that projecting these vectors onto the vocabulary space resulted in high logit values for the tokens corresponding to the words meant to be excluded. In other words, adding these vectors to the model’s residual stream actually increased the probability of generating the very keywords we intended to exclude. To address this issue, instead of computing exclusion vectors, we compute the vectors for inclusion and then subtract them from the model’s residual stream. This technique effectively steers the model away from using the words captured by the vector. Figure 6b shows the results of steering for keyword exclusion using this approach. We observe that subtracting the inclusion vectors significantly reduces the frequency of the undesired keywords in the model’s responses, both in cases where no textual instructions are present and where explicit exclusion instructions are provided in the input. § 5 reports an example where the instruction to exclude a specific word is ineffective on its own, but applying steering successfully enforces the constraint.
@@ 6 Multi-instruction Steering
Figure 7 shows results from experiments with steering two instructions at the same time: format (all 13 instructions) and length (Figure 7a), as well as lowercase and keyword exclusion (Figure 7b). In both cases, steering leads to improvements on both axes. These findings suggest that steering for multiple instructions at once is feasible and can lead to performance gains across different constraints. However, we anticipate that issues may arise in certain cases, particularly when dealing with conflicting instructions, where refinement may be necessary to balance competing constraints.
@@ 7 Cross-model Steering
Figure 8: Cross-model Steering. (a) Without text instructions, cross-model steering improves constraint satisfaction more than same-model steering. (b) With text instructions, cross-model steering increases accuracy for Gemma 2 2B, but not for Gemma 2 9B. (c) Cross-model steering for length instructions shortens output as steering weight increases.
Instruction-tuned models show improved instruction-following abilities, suggesting that they may also be better at constructing meaningful representations of instructions. In this context, we ask: can we leverage the representations computed on instruction-tuned models to steer a base model more effectively? This idea is motivated by prior work showing that the weights of language models do not change significantly after instruction tuning (Lee et al., 2024; Jain et al., 2024). Additionally, linear representations derived through mean activation difference and sparse autoencoders (Huben et al., 2024) have been shown to be transferable between base and chat models (Panickssery et al., 2024; Kissane et al., 2024). To study this in our setting, we conduct experiments with the base and instruction-tuned versions of Gemma 2 2B and 9B. We focus on format instructions, using the same data and procedures outlined in § 3. The key difference is that we apply steering vectors computed on the instruction-tuned models to steer the base models, comparing this approach to same-model steering (vectors computed on the base model).
When inputs consist of base queries only (Figure 8a), steering the base Gemma models with same-model vectors results in only modest improvements. However, when using vectors from instruction-tuned counterparts, the base models satisfy the constraints much more frequently, indicating the transferability of the instruction representations. When explicit input instructions are introduced (Figure 8b), Gemma 2B sees further accuracy gains, with cross-model steering outperforming same-model steering (40.5% vs. 36.2%). For Gemma 9B, however, steering yields no significant improvement, suggesting that the performance bottleneck might not be the quality of the instruction representation. We also explore cross-model steering for length instructions. As described in § 4, we compute a steering vector for conciseness from the instruction-tuned version of Gemma 9B and apply it to the base version. Figure 8c shows that increasing the steering weight shortens the outputs, demonstrating that representations learned in instruction-tuned models can still meaningfully modulate the output in the base model’s space.
@@ 8 Related Work
Language Model Representations. Our approach is inspired by prior research that shows it is possible to obtain vector representations encoding information about tasks on which a language model has been trained (Ilharco et al., 2023; Huang et al., 2024) or tasks learned in context (Hendel et al., 2023; Todd et al., 2024). These studies are part of a broader body of work that examines the linear representation of features such as truthfulness (Li et al., 2023; Azaria & Mitchell, 2023; Marks & Tegmark, 2024), sentiment (Tigges et al., 2024), harmlessness (Zou et al., 2023; Zheng et al., 2024), sycophancy (Perez et al., 2023; Panickssery et al., 2024; Sharma et al., 2024), factual knowledge (Gurnee & Tegmark, 2024), and refusal (Arditi et al., 2024). In addition, recent works have employed sparse autoencoders to identify feature directions in an unsupervised manner (Bricken et al., 2023; Huben et al., 2024; Templeton et al., 2024). A shared hypothesis across these works is that LLMs represent features or concepts as linear directions in activation space (Mikolov et al., 2013; Bolukbasi et al., 2016; Elhage et al., 2021; Nanda et al., 2023; Park et al., 2024; Olah, 2024). While recent studies suggest that not all features may be linearly encoded (Engels et al., 2024; Csordás et al., 2024), the linearity assumption has been effective for both concept erasure (Ravfogel et al., 2020; Belrose et al., 2023; Shao et al., 2023; Guerner et al., 2024) and model steering.
@@ 9 Conclusion
We demonstrated the effectiveness of activation steering for improving language models’ adherence to instructions. By computing steering vectors based on activation differences between inputs with and without instructions, we guide models to follow constraints related to format, length, and word inclusion/exclusion. These vectors capture meaningful instruction representations, enabling models to meet constraints even without explicit instructions, while also enhancing performance when instructions are present. Additionally, we show that models can handle multiple constraints simultaneously, offering a flexible framework for controlled language generation. Finally, we explore cross-model steering, revealing that vectors computed on instruction-tuned models can improve the behavior of base models, often surpassing same-model steering.
@@
In § 2.2, we outline our methodology in detail. Information about the datasets we used, as well as the procedures for obtaining and augmenting them, is provided in § 2.3, along with Apps. B and C. § 2.3 also contains the details of the evaluation metrics we employed and the process for generating model outputs. Appendix E provides further details on the implementation of our quality score metric. Additional implementation and experimental details are included in Appendix F.
@@ Acknowledgments
Appendix A List of Instructions
@@
Appendix B Extracting IFEval Base Queries
To obtain base queries without instructions, we use GPT-4o (OpenAI, 2024) [^6] https://openai.com/index/hello-gpt-4o/ to strip the instructions from the base queries, as the IFEval dataset does not explicitly annotate the instructions within the prompts. We use a one-shot prompt a fixed in-context example. The prompt used is reported in Table 5. These base queries are then used to generate additional base query-instruction combinations, augmenting the original dataset. Further details on the data used for steering vector computation and evaluation are provided in Appendix C.
@@ Appendix B Extracting IFEval Base Queries
Appendix C Data and Vector Computation Details
@@
Format Instructions. The steering vectors and layer selection for format instructions are computed using a separate set of synthetically generated prompts. To generate a synthetic dataset containing all base prompts with all instructions from the IFEval dataset, we start with the base prompts as described in Appendix B. We then create an augmented version of the IFEval dataset by combining each base prompt with every available instruction. This is done by prompting GPT-4o using in-context examples, as shown in Table 6. When adding a constraint to a base prompt, we include a randomly selected in-context example of the same type of constraint from the single-constraint dataset. These single-constraint examples are generated by prompting GPT-4o to remove all but one instruction from each prompt. For this, we use a fixed, manually curated in-context example. This procedure is carried for each base query and each format instruction. We then filter out any base query and instruction combinations that appear in the evaluation set to ensure that the base queries used for steering vector computation do not overlap with those in the evaluation prompts. This process yields approximately 450 examples for each instruction, which are used for both the steering vector computation and for validation. We provide examples of data generated using this procedure in Table 7.
@@
For format instruction evaluation, we utilize the format-related subset of IFEval queries (163 in total), which includes the 13 distinct instructions detailed in Appendix A.
@@
Length Instructions. For length instructions, the steering vectors are computed using a fixed set of 50 IFEval base queries, obtained as described in Appendix B. To this set, we append various manually annotated phrasings of instructions that request the model to generate concise responses, such as “Be concise” or “The answer should be brief.”
@@
The evaluation data used in § 4 consists of a separate set of 200 base queries. Although IFEval contains some length-related instructions, it offers limited examples with significant variation in length constraints and expressed in different ways (e.g., “I don’t want anything longer than 30 words,” or “Answer using five paragraphs”). We opt for synthetically generating data to enable consistent evaluation across a set of constraints varying over a narrower range (1 to 5 sentences).
@@
Word-specific Instructions. For word-specific instructions, we compute vector representations for inclusion constraints, where the model is instructed to include a specific keyword in the output (e.g., “The output should contain the word {$w$}” or “Ensure that the word {$w$} is included in the response”). For each keyword, we use a set of 20 base queries randomly sampled from those generated as described in Appendix B, ensuring none of the queries used for vector computation are reused in evaluation. The topics and content of these base queries are often semantically unrelated to the keyword being included.
@@
For evaluation, we use IFEval examples containing keyword inclusion and exclusion instructions. Many examples contain multiple keyword constraints in a single query, so we separate these into individual prompts, each requesting the inclusion or exclusion of a single keyword. This process yields 86 evaluation prompts for keyword inclusion and 117 for keyword exclusion.
@@
Capital Word Freq. 18 20 11 - 12 24 16 8
English Capital 28 18 11 11 18 30 28 8
English Lowercase 18 30 17 7 28 6 22 -
@@
Json Format 16 6 - 13 16 12 - 10
Multiple Sections - 6 - 9 - 24 - 28
@@
Highlight Text 26 18 5 - 26 26 12 10
Title - - - - - - - -
No Comma 10 12 11 11 14 26 22 -
@@
Language Fa 20 - - - 18 20 22 -
Language Fi 18 - 15 - 16 8 22 -
Language Gu - - - - - - - -
Language Hi 18 - 15 7 26 - 22 12
Language It 16 - 15 - 16 - 20 -
Language Mr 24 - 17 15 18 18 24 -
Language Pa - - - - 22 - - 8
@@
Language Ru 16 - 15 - 16 - 22 -
Language Sw 20 - - - - 14 22 20
Language Ta 18 - - - 18 - 22 -
Language Te 24 - 19 - 16 - 20 -
Language Th 18 - 15 - 18 - 20 -
Language Ur 24 - - 11 - - 28 -
Language Vi 20 - 15 - 18 - 20 -
Appendix D Steering Layer Selection
In Table 8, we provide the layer indices used in our steering procedure for each instruction in the format subset (excluding language) across all models and evaluation settings (with and without explicit text instructions). Layer selection is based on validation instruction-following accuracy, evaluated on a set of 96 held-out examples (8 per instruction), sampled from the synthetically generated prompts described in Appendix C. We compare the validation scores for each layer against a baseline without steering to ensure that the steering intervention results in an improvement. If no layer shows an improvement, steering is not applied at test time. In the event of a tie between layers with the highest score, the earliest layer is chosen.

### v2

@@ 5 Word-specific Instructions
Word Inclusion. For this set of instructions, we compute word-specific steering vectors. To generate these vectors, we append different phrasings of a request to include a specific word $w$ (e.g., “please include the word {$w$} in your response”) to a set of base queries. These prompts are used to compute a vector representation for the “include word {$w$}” instruction. While this requires a separate steering vector for each keyword at inference time, the vectors can be generated on-the-fly using arbitrary base queries unrelated to the keyword itself. Furthermore, these vectors can be computed with a small number of examples; in our experiments, we use only 20 examples. The models are then evaluated on the subset of IFEval that contains keyword inclusion/exclusion constraints. [^5] IFEval prompts may contain instructions for the inclusion or exclusion of multiple keywords in the same example. We separate such cases into inputs with single-keyword instructions.
@@
Word Exclusion. For keyword exclusion, we initially used the same procedure, appending instructions like “ensure the word $w$ does not appear in the answer” to base queries and computing the corresponding steering vectors. However, upon inspecting the steering vectors, we found that projecting these vectors onto the vocabulary space resulted in high logit values for the tokens corresponding to the words meant to be excluded. In other words, adding these vectors to the model’s residual stream actually increased the probability of generating the very keywords we intended to exclude. To address this issue, instead of computing exclusion vectors, we compute the vectors for inclusion and then subtract them from the model’s residual stream. This technique effectively steers the model away from using the words captured by the vector. Figure 6b shows the results of steering for keyword exclusion using this approach. We observe that subtracting the inclusion vectors significantly reduces the frequency of the undesired keywords in the model’s responses, both in cases where no textual instructions are present and where explicit exclusion instructions are provided in the input. § 5 reports an example where the instruction to exclude a specific word is ineffective on its own, but applying steering successfully enforces the constraint.
@@ 6 Multi-instruction Steering
Figure 7 shows results from experiments with steering two instructions the same time: format (all 13 instructions) and length (Figure 7a), as well as lowercase and keyword exclusion (Figure 7b). In both cases, steering leads to improvements on both axes. These findings suggest that steering for multiple instructions at once is feasible and can lead to performance gains across different constraints. However, we anticipate that issues may arise in certain cases, particularly when dealing with conflicting instructions, where further refinement may be necessary to balance competing constraints.
@@ 7 Cross-model Steering
Figure 8: Cross-model Steering. (a,b) Steering vectors from instruction-tuned models transfer effectively to base models, with cross-model steering outperforming same-model steering in Gemma 2 2B. (c) Cross-model steering for length instructions shortens output as steering weight increases.
Instruction-tuned models show improved instruction-following abilities, suggesting they may also form more effective instruction representations. This raises the question: can we leverage the representations computed on instruction-tuned models to steer a base model more effectively? This idea is motivated by prior work showing that fine-tuning reinforces existing mechanisms in language models (Prakash et al., 2024) and that instruction tuning causes minimal weight changes (Lee et al., 2024; Jain et al., 2024). Additionally, linear representations derived through mean activation difference and sparse autoencoders (Huben et al., 2024) have been shown to be transferable between base and chat models (Panickssery et al., 2024; Kissane et al., 2024). To study this in our setting, we conduct experiments with the base and instruction-tuned versions of Gemma 2 2B and 9B. We focus on format instructions, using the same data and procedures outlined in § 3. The key difference is that we apply steering vectors computed on the instruction-tuned models to steer the base models, comparing this approach to same-model steering (vectors computed on the base model).
In both the base query-only setting (Figure 8a) and with explicit instructions (Figure 8b), steering base models using vectors computed on instruction-tuned counterparts yields performance gains that demonstrate the transferability of instruction representations. Notably, in the case of Gemma 2 2B, cross-model steering outperforms same-model steering across both settings, highlighting its potential for more effective instruction adherence in smaller models. We also explore cross-model steering for length instructions. As described in § 4, we compute a steering vector for conciseness from the instruction-tuned version of Gemma 9B and apply it to the base version. Figure 8c shows that increasing the steering weight shortens the outputs, demonstrating that representations learned in instruction-tuned models can still meaningfully modulate the output in the base model’s space.
@@ 8 Related Work
Language Model Representations. Our approach is inspired by prior research that shows it is possible to obtain vector representations encoding information about tasks on which a language model has been trained (Ilharco et al., 2023; Huang et al., 2024) or tasks learned in context (Hendel et al., 2023; Todd et al., 2024). These studies are part of a broader body of work that examines the linear representation of features such as truthfulness (Li et al., 2023; Azaria & Mitchell, 2023; Marks & Tegmark, 2024), sentiment (Tigges et al., 2024), harmlessness (Zou et al., 2023; Zheng et al., 2024), sychophancy (Perez et al., 2023; Panickssery et al., 2024; Sharma et al., 2024), factual knowledge (Gurnee & Tegmark, 2024), and refusal (Arditi et al., 2024). In addition, recent works have employed sparse autoencoders to identify feature directions in an unsupervised manner (Bricken et al., 2023; Huben et al., 2024; Templeton et al., 2024). A shared hypothesis across these works is that LLMs represent features or concepts as linear directions in activation space (Mikolov et al., 2013; Bolukbasi et al., 2016; Elhage et al., 2021; Nanda et al., 2023; Park et al., 2024; Olah, 2024). While recent studies suggest that not all features may be linearly encoded (Engels et al., 2025; Csordás et al., 2024), the linearity assumption has been effective for both concept erasure (Ravfogel et al., 2020; Belrose et al., 2023; Shao et al., 2023; Guerner et al., 2024) and model steering.
@@ 9 Conclusion
We demonstrated the effectiveness of activation steering for improving language models’ adherence to instructions. By computing steering vectors based on activation differences between inputs with and without instructions, we guide models to follow constraints related to format, length, and word inclusion/exclusion. These vectors capture meaningful instruction representations, enabling models to meet constraints even without explicit instructions, while also enhancing performance when instructions are present. Additionally, we show that models can handle multiple constraints simultaneously, offering a flexible framework for controlled language generation. Finally, we explore cross-model steering, revealing that vectors computed on instruction-tuned models can improve the behavior of base models, in some cases surpassing same-model steering.
@@
We submitted our code and data as supplementary materials, and they will be open-sourced upon publication. § 2.2 outlines our methodology in detail. Information about the datasets we used, as well as the procedures for obtaining and augmenting them, is provided in § 2.3, along with Apps. C and D. § 2.3 also contains the details of the evaluation metrics we employed and the process for generating model outputs. Appendix F provides further details on the implementation of our quality score metric. Additional implementation and experimental details are included in Appendix G.
@@ Acknowledgments
Appendix A Limitations and Future Work
While our work demonstrates the potential of activation steering for improving instruction adherence, it has multiple limitations that highlight avenues for future research. Activation steering involves several degrees of freedom, including the selection of the model layer to intervene at, determining the steering weight, and identifying where in the input or output sequence to apply the intervention. While our method systematically selects these parameters based on paired examples that differ by the presence of an instruction, it does not fully explore the space of possible interventions, leaving room for optimization. For instance, the quality of instruction representations directly impacts the effectiveness of steering. Representations computed from inputs where the model fails to adhere to instructions may reduce the efficacy of steering. Future work could explore filtering out such representations during vector computation to improve results, although this would require additional instruction-following accuracy checks, which our current approach avoids. Furthermore, our method applies a fixed steering weight throughout the generation process. Dynamically modulating the steering weight during generation (Scalena et al., 2024) could adapt the intervention to the evolving context and improve performance.
Another area for exploration is the effect of the number of examples used to compute steering vectors. Previous research indicates that clean concept representations can be obtained with relatively few examples (e.g., 128; Arditi et al., 2024). In our experiments, steering vectors computed from as few as 20 examples proved effective. However, we did not extensively investigate how varying the size of the example set affects the quality of these vectors. Future research could focus on optimizing this balance between computational efficiency and steering performance.
While we consider a reasonably large and diverse set of instructions in our experiments, the range of possible real-world user requests is vast and cannot be exhaustively represented. Additionally, our use of perplexity (computed with GPT-2; Radford et al., 2019) as a proxy for output quality focuses on detecting issues in fluency and coherence but may miss subtler forms of degradation, such as factual inaccuracies or logical inconsistencies. Incorporating more nuanced quality metrics or task-specific evaluations could improve parameter selection.
Finally, an interesting direction for future work stems from our cross-model steering experiments. In our work, we experiment with transferring steering vectors from instruction-tuned to base models of the same family and parameter size. An interesting avenue for future research is exploring the transferability of steering vectors across models with different architectures or sizes, potentially by learning transformations that map the latent space of one model to another.
Appendix B List of Instructions
@@
Appendix C Extracting IFEval Base Queries
To obtain base queries without instructions, we use GPT-4o (OpenAI, 2024) [^6] https://openai.com/index/hello-gpt-4o/ to strip the instructions from the base queries, as the IFEval dataset does not explicitly annotate the instructions within the prompts. We use a one-shot prompt a fixed in-context example. The prompt used is reported in Table 5. These base queries are then used to generate additional base query-instruction combinations, augmenting the original dataset. Further details on the data used for steering vector computation and evaluation are provided in Appendix D.
@@ Appendix C Extracting IFEval Base Queries
Appendix D Data and Vector Computation Details
@@
Format Instructions. The steering vectors and layer selection for format instructions are computed using a separate set of synthetically generated prompts. To generate a synthetic dataset containing all base prompts with all instructions from the IFEval dataset, we start with the base prompts as described in Appendix C. We then create an augmented version of the IFEval dataset by combining each base prompt with every available instruction. This is done by prompting GPT-4o using in-context examples, as shown in Table 6. When adding a constraint to a base prompt, we include a randomly selected in-context example of the same type of constraint from the single-constraint dataset. These single-constraint examples are generated by prompting GPT-4o to remove all but one instruction from each prompt. For this, we use a fixed, manually curated in-context example. This procedure is carried for each base query and each format instruction. We then filter out any base query and instruction combinations that appear in the evaluation set to ensure that the base queries used for steering vector computation and layer selection do not overlap with those in the evaluation prompts. This process yields approximately 450 examples for each instruction, which are used for both the steering vector computation and for validation. We provide examples of data generated using this procedure in Table 7.
@@
For validation, we use a set of 96 examples (8 per instruction), sampled from the synthetically data described above. For format instruction evaluation, we utilize the format-related subset of IFEval queries (163 in total), which includes the 13 distinct instructions detailed in Appendix B.
@@
Length Instructions. For length instructions, the steering vectors are computed using a fixed set of 50 IFEval base queries, obtained as described in Appendix C. To this set, we append various manually-annotated phrasings of instructions that request the model to generate concise responses, such as “Be concise” or “The answer should be brief.”
@@
The evaluation data used in § 4 and Appendix J consists of a separate set of 200 base queries. Although IFEval contains some length-related instructions, it offers limited examples with significant variation in length constraints and expressed in different ways (e.g., “I don’t want anything longer than 30 words,” or “Answer using five paragraphs”). We opt for synthetically generating data to enable consistent evaluation across a set of constraints varying over a narrower range (1 to 5 sentences).
@@
Word-specific Instructions. For word-specific instructions, we compute vector representations for inclusion constraints, where the model is instructed to include a specific keyword in the output (e.g., “The output should contain the word {$w$}” or “Ensure that the word {$w$} is included in the response”). For each keyword, we use a set of 20 base queries randomly sampled from those generated as described in Appendix C, ensuring none of the queries used for vector computation are reused in evaluation. The topics and content of these base queries are often semantically unrelated to the keyword being included. The same procedure is carried out for word-exclusion instructions in Appendix K.
@@
For validation, we use GPT-4o to synthetically generate a set of questions similar to the base queries in IFEval. Additionally, the prompt requests the generation of a list of words likely to appear in the answer to each question. These question-word pairs (276 in total) are used for grid search validation of word inclusion and exclusion. For evaluation, we use IFEval examples containing keyword inclusion and exclusion instructions. Many examples contain multiple keyword constraints in a single query, so we separate these into individual prompts, each requesting the inclusion or exclusion of a single keyword. This process yields 86 evaluation prompts for keyword inclusion and 117 for keyword exclusion.
@@
Capital Word Freq. 18 20 21 23 12 24 16 -
Capitalize 28 22 11 11 28 30 28 8
Lowercase 18 30 17 7 18 6 22 -
@@
JSON Format - 6 - 13 16 - - 10
Multiple Sections - 8 - 9 - - - 28
@@
Highlighted Text 26 18 5 - 24 26 12 10
Title - - - - - - 34 -
No Comma 10 12 11 5 14 26 22 -
@@
Language Fa 20 - - - - - 22 -
Language Fi 20 - 15 - 20 12 22 -
Language Gu - 14 - - - - - -
Language Hi 22 14 15 7 - - 22 12
Language It 16 - 15 - 18 - 20 -
Language Mr 30 14 17 15 - - 24 -
Language Pa - - - - - - - 8
@@
Language Ru 16 - 15 - 18 - 22 -
Language Sw 20 - - - - 24 22 20
Language Ta - 6 - - - - 22 -
Language Te - - - 7 - - 20 -
Language Th - - - 5 - - 20 -
Language Ur 28 - - 15 - - 28 -
Language Vi 20 - 15 - - 12 20 -
Appendix E Steering Layer and Weight Selection
When selecting the steering layer and weight, our goal is to identify a combination that improves instruction-following accuracy without compromising the fluency or comprehensiveness of the output. While instruction-following accuracy can be quantified through specific checks, evaluating fluency and comprehensiveness is more challenging, as thorough evaluations, such as our GPT-4o-based procedure, are computationally expensive and impractical for validation. To address this, we use perplexity computed by a smaller model as a proxy (GPT-2; Radford et al., 2019). The rationale is that very low perplexity values can indicate outputs where steering has significantly degraded quality. For each layer-weight combination, we compute the fraction of low-perplexity outputs (below a threshold) and exclude configurations where this fraction is non-zero. [^7] In non-instruction-tuned models low perplexity may occur even without steering. In these cases we exclude layers if the post-steering fraction exceeds the baseline value observed before steering. Among the remaining configurations, we select the one with the highest validation accuracy. Additionally, we compare the validation accuracy for each layer against a baseline without steering to ensure the intervention provides an improvement. If no layer improves upon the baseline, steering is not applied at test time. In cases where multiple layers achieve the same highest accuracy, we select the earliest layer. The perplexity threshold provides a mechanism for balancing instruction-following accuracy against the risk of quality degradation. A higher threshold results in a more conservative steering procedure but may exclude valid parameter combinations that could yield higher accuracy. In our experiments, we set the threshold to 2.5 for format instructions and 4 for word-specific instructions. This systematic approach ensures effective steering while maintaining output quality.